# LCQ: Low-Rank Codebook based Quantization for Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs), such as GPT [3], GLM [37], LLaMA [29] and LLava [18], have recently demonstrated promising performance in many tasks including both understanding and generation of natural language, image, video and speech. The number of parameters in LLMs is typically very large. For example, the number of parameters in GPT [3] can reach up to 175 billion. A large number of parameters will result in high storage and computational cost, which has become a challenge for deploying LLMs in real applications.
@@
Weight quantization [17, 9, 6, 34], which represents (quantizes) each high bit-width weight by a low bit-width value, has been widely used for model compression. Weight quantization can reduce both storage and computational cost for model inference [9], and hence can facilitate model deployment in real applications. Naive weight quantization will typically cause accuracy degradation, especially for cases when the compression ratio is high, i.e., the bit-width after quantization is too low. Existing weight quantization methods for LLMs have proposed different techniques to mitigate the accuracy degradation caused by quantization. For example, GPTQ [9] uses output reconstruction error rather than simple parameter reconstruction error for quantization. GPTQ [9] primarily proposes a learning method to compensate for quantization residuals and does not perform additional optimizations for the quantization codebook. AWQ [17] and OmniQuant [28] mainly focus on learning the quantization codebook, further reducing the performance loss of the quantized model. Although existing weight quantization methods can achieve accuracy comparable to the original model with 4-bit or higher-bit quantization, there remains a significant accuracy degradation for quantization when the bit-width is lower than 4.
@@
One common characteristic of most existing weight quantization methods for LLMs is that they use a rank-one codebook for quantization. Specifically, the quantization codebook for weights can be seen as the outer product of a scaling vector [17, 9, 6] and a quantization point set (QPS) vector [24]. Each element in the scaling vector corresponds to a scaling factor for a group of weights and the QPS vector is constructed from quantiles of uniform distribution  [9, 17] or Gaussian distribution [6]. Although the rank-one codebook can reduce the storage cost of the model, it has limited representation ability. This might be one of the main reasons to explain the significant accuracy degradation for existing quantization methods when the bit-width is lower than 4.
@@
In this paper, we propose a novel weight quantization method, called low-rank codebook based quantization (LCQ), for LLMs. The main contributions of LCQ are outlined as follows:
@@ 2 Related Work
Existing weight quantization methods for LLMs typically employ post-training quantization (PTQ) which is relatively fast and does not require access to the full original training data [9, 17, 6]. PTQ methods such as BRECQ [16], OBC [8] and QDrop [31] are designed for quantizing traditional non-LLMs, which cannot be directly used for LLMs due to their high optimization cost. Recently, some PTQ methods have been specially designed for LLM quantization. GPTQ [9] primarily proposes a learning method to compensate for quantization residuals and uses output reconstruction error as the objective function. However, GPTQ does not perform optimization for the quantization codebook, which results in significant accuracy loss with low bit-width. AWQ [17] and OmniQuant [28] mainly focus on learning the quantization codebook, and the quantization indices are directly computed based on the index of the quantization value that is nearest to the full-precision weight. The codebooks of AWQ and OmniQuant are rank-one, limiting the representation ability. Very recently, there appear some PTQ methods like OWQ [14], SpQR [7] and SquzzeLLM [13], which mainly address the issue of outliers in activations by assigning higher bit-widths to the channels/connections of activation outliers. Our LCQ is orthogonal to methods like OWQ, SpQR and SqueezeLLM, and can be further integrated with these methods.
Moreover, some LLMs compression methods [5, 33, 36, 32] integrate weight quantization and activation quantization for further accelerating LLMs inference. However, the accuracy drops significantly when quantizing both weights and activations.
PTQ methods have also been integrated with parameter-efficient fine tuning (PEFT). Representative methods include QLoRA [6], PEQA [12], QA-LoRA [34], and LoftQ [15]. But the main goal of these methods is for PEFT rather than directly quantiting the original LLMs. Our LCQ can also be used for PEFT, but this is not the focus of this paper.
@@ 3 Methodology
As in existing weight quantization methods for LLMs [9, 17, 28], we also employ post-training quantization in our LCQ. In post-training quantization, we have a calibration dataset, which is denoted as $\{\mathbf{X}_{i}\}_{i=1}^{N_{X}}$. Here, each calibration sample $\mathbf{X}_{i}\in\mathbb{R}^{L\times D}$ has a sequence length of $L$ and feature dimension of $D$, and $N_{X}$ is the total number of calibration samples. As in AWQ [17] and OmniQuant [28], we perform grouped quantization for the weights of linear layers. More specifically, for a linear layer with $N_{W}$ weights, we first divide all the weights into $N_{V}$ subsets, with each subset having $N_{W}/N_{V}$ weights. Then for each subset of weights, we group its weights into $N_{G}$ groups. We denote each subset of weights as $\mathbf{W}\in\mathbb{R}^{N_{G}\times G}$, in which $G$ denotes the weight group size and $N_{G}=\frac{N_{W}}{N_{V}\times G}$ denotes the number of weight groups per subset. A quantization codebook containing quantization values is maintained for each group of weights. For weight quantization with $b$ bits, the number of all quantization values of each group of weights is $N_{Q}=2^{b}$. We denote the quantization codebook of all groups as $\mathbf{C}\in\mathbb{R}^{N_{G}\times N_{Q}}$, the $i$th row of which is the codebook corresponding to the $i$th group of weights. The quantization function operates on the weights in an element-wise way, which quantizes each weight to the nearest quantization value in the codebook. Let $\mathbf{Z}\in[0,1,\cdots,N_{Q}]^{N_{G}\times G}$ denote the quantization index of $\mathbf{W}$, in which $Z_{i,j}$ denotes the quantization index of $W_{i,j}$. We have
@@
Then we can use the quantized weight matrix $\mathbf{W}^{Q}$ for inference. In real deployment, we only need to store the codebook matrix $\mathbf{C}$ and the quantization index matrix $\mathbf{Z}$ [11]. From (1), we can find that $\mathbf{Z}$ is computed based on $\mathbf{C}$. Hence, $\mathbf{C}$ is the key for quantization, which will affect accuracy, storage cost and computational cost.
@@ 3.1 Low-Rank Codebook
Figure 1: Illustration of low-rank codebook based quantization.
In AWQ [17] and GPTQ [9], the intervals between quantization values are equal, and hence it is not necessary to explicitly store the quantization codebook $\mathbf{C}$. More specifically, $\mathbf{C}$ can be represented by two vectors, a scaling vector $\mathbf{S}_{1}\in\mathbb{R}^{1\times N_{G}}$ and a fixed quantization point set (QPS) vector $\mathbf{V}_{1}\in\mathbb{R}^{1\times N_{Q}}$. $\mathbf{V}_{1}=[-1,-1+\frac{2}{2^{b}-1},\cdots,1-\frac{2}{2^{b}-1},1]$ is a uniformly spaced vector between $[-1,1]$. $\mathbf{S}_{1}$ can be adaptively obtained based on the weight matrix $\mathbf{W}$. The codebook is then obtained by computing the outer product of $\mathbf{S}_{1}$ and $\mathbf{V}_{1}$. Furthermore, to adapt to asymmetric weight distribution, a quantization offset $\mathbf{B}\in\mathbb{R}^{N_{G}\times 1}$ is introduced, with each group of weights having its own quantization offset. The final quantization codebook can be formulated as $\mathbf{C}=\mathbf{S}^{T}_{1}\mathbf{V}_{1}-\mathbf{B}$, which is a rank-one matrix. The storage cost for $\mathbf{S}_{1}$, $\mathbf{V}_{1}$ and $\mathbf{B}$ is $O(2*N_{G}+N_{Q})$, which is much lower than $O(N_{G}*N_{Q})$ for storing $\mathbf{C}$. Although rank-one codebook can reduce the storage cost, it has limited representation ability.
@@
When $N_{D}>1$, the rank of the codebook $\mathbf{C}$ can be larger than one. When $N_{D}=1$, the codebook $\mathbf{C}$ degenerates to rank-one codebook as in existing methods [17, 9]. Hence, our LCQ has a stronger representation ability than existing methods, including existing rank-one codebook based methods as special cases. We will show that LCQ with $N_{D}>1$ will achieve better accuracy than LCQ with $N_{D}=1$ in Section 4. This verifies the stronger representation ability of our low-rank codebook compared to the rank-one codebook. Figure 1 shows the low-rank codebook with $N_{D}=2$.
@@ 3.2 Objective Function
As illustrated in Figure 1, LCQ starts with a pre-trained model and tries to find the optimal $\mathbf{S}$, $\mathbf{V}$ and $\mathbf{B}$ for minimizing the loss caused by quantization. We use the output reconstruction error of a Transformer block as the objective function like the methods in [1, 16, 35], which can make the parameters of a Transformer block be optimized simultaneously. Compared to the methods that use the output reconstruction error of a single linear layer as the objective function [9, 17], using the output reconstruction error of a Transformer block can avoid falling into local optima and prevent the accumulation of quantization error. A Transformer block in the Transformer-based models [30] typically contains the following linear layers: the query/key/value linear projection layers $\mathbf{W}^{qproj},\mathbf{W}^{kproj},\mathbf{W}^{vproj}$, the output projection layer $\mathbf{W}^{oproj}$ of the multi-head self-attention, and two linear layers $\mathbf{W}^{fc1}$ and $\mathbf{W}^{fc2}$ of the feed forward network. Therefore, the parameters of a Transformer block can be represented as $\mathcal{W}=\left[\mathbf{W}^{qproj},\mathbf{W}^{kproj},\mathbf{W}^{vproj},\mathbf{W}^{oproj},\mathbf{W}^{fc1},\mathbf{W}^{fc2}\right]$. Similarly, parameters $\mathcal{S},\mathcal{V},\mathcal{B}$ are used to represent all $\mathbf{S}$s, $\mathbf{V}$s, $\mathbf{B}$s within a Transformer block. We define the objective function to minimize the norm of the difference between the output with quantized weights and the output with unquantized weights in a Transformer block as follows:
$\min_{\mathcal{S},\mathcal{V},\mathcal{B}}$ $\sum_{i=1}^{N_{X}}||h(\tilde{\mathbf{X}}_{i},\mathcal{W}^{Q})-h(\mathbf{X}^{fp}_{i},\mathcal{W})||_{F}^{2}+||h(\tilde{\mathbf{X}}_{i},\mathcal{W}^{Q})-h(\tilde{\mathbf{X}}_{i},\mathcal{W})||_{F}^{2}.$ (4)
where $h(\cdot)$ denotes the function of a Transformer block, and $N_{X}$ is the number of calibration samples. $\mathbf{X}^{fp}_{i}$ represents the feature with all previous blocks full precision and $\tilde{\mathbf{X}}_{i}$ represents the feature with all previous blocks quantized. $\mathcal{W}^{Q}=Q(\mathcal{W},\Phi(\mathcal{S},\mathcal{V},\mathcal{B}))$ represents the quantized weights in a Transformer block. $h(\tilde{\mathbf{X}}_{i},\mathcal{W}^{Q})$ represents the output feature of the Transformer block with all blocks quantized including this block. The objective function contains two items, one use $h(\mathbf{X}^{fp}_{i},\mathcal{W})$ as the reconstruction targets and another uses $h(\tilde{\mathbf{X}}_{i},\mathcal{W})$. Note that the quantization algorithm does not change the model weights $\mathbf{W}$ [28], in order to avoid forgetting existing knowledge. By fixing $\mathbf{W}$, we can also reduce the optimization overhead by reducing the number of parameters to be optimized.
@@ 3.3.1 Gradient Approximation for Quantization Function
The $argmin(\cdot)$ operation in (1) will make the gradient unable to backpropagate through this function. This makes it challenging to learn the quantization codebook. We propose a gradient approximation strategy to address this issue.
As in N2UQ [19], we rewrite the quantization function into a form that adds up multiple segments, each corresponding to the interval between two adjacent quantization values. The form is as follows:
$Q(W_{i,j},\mathbf{C}_{i})$ $=C_{i,\pi_{i}(1)}+\sum_{k=1}^{N_{Q}-1}\left(C_{i,\pi_{i}(k+1)}-C_{i,\pi_{i}(k)}\right)*\chi_{0.5}^{(k)}\left(\text{clip}\left(\frac{W_{i,j}-C_{i,\pi_{i}(k)}}{C_{i,\pi_{i}(k+1)}-C_{i,\pi_{i}(k)}},0,1\right)\right),$ (5)

### v2

@@ 1 Introduction
Large language models (LLMs), such as GPT (Brown et al., 2020), GLM (Zeng et al., 2023), LLaMA (Touvron et al., 2023) and LLaVA (Liu et al., 2022a), have recently demonstrated promising performance in many tasks including both understanding and generation of natural language, image, video and speech. However, as these models grow in size, their parameter counts also increase dramatically. For instance, the number of parameters in GPT (Brown et al., 2020) can reach up to 1.8 trillion. A large number of parameters will result in high storage and computational cost, which has become a challenge for deploying LLMs in resource-constrained devices such as mobile phones or edge devices.
Weight quantization (Lin et al., 2024; Frantar et al., 2023; Dettmers et al., 2023; Xu et al., 2024), which represents (quantizes) each high bit-width weight by a low bit-width value, has been widely used for model compression. Weight quantization can reduce both storage and computational cost for model inference (Frantar et al., 2023), and hence can facilitate model deployment in real applications. For example, quantizing a model from 16-bit floating-point precision to 4-bit integers can reduce the model size by a factor of four, enabling deployment on devices with limited memory.
@@
Despite the benefits, naive weight quantization will typically cause accuracy degradation, especially for cases when the compression ratio is high, i.e., the bit-width after quantization is too low. Existing weight quantization methods for LLMs have proposed different techniques to mitigate the accuracy degradation caused by quantization. For example, GPTQ (Frantar et al., 2023) uses output reconstruction error rather than simple parameter reconstruction error for quantization. GPTQ (Frantar et al., 2023) primarily proposes a learning method to compensate for quantization residuals and does not perform additional optimizations for the quantization codebook. AWQ (Lin et al., 2024) and OmniQuant (Shao et al., 2024) mainly focus on learning the quantization codebook, further reducing the performance loss of the quantized model. Although existing weight quantization methods can achieve accuracy comparable to the original model with 4-bit or larger bit-width quantization, there remains a significant accuracy degradation for quantization when the bit-width is lower than 4.
@@
One common characteristic of most existing weight quantization methods for LLMs is that they use a rank-one codebook for quantization. Specifically, the quantization codebook for weights can be seen as the outer product of a scaling vector (Lin et al., 2024; Frantar et al., 2023; Dettmers et al., 2023) and a quantization point set (QPS) vector (Oh et al., 2022). Each element in the scaling vector corresponds to a scaling factor for a group of weights and the QPS vector is constructed from quantiles of uniform distribution  (Frantar et al., 2023; Lin et al., 2024) or Gaussian distribution (Dettmers et al., 2023). Although the rank-one codebook can reduce the storage cost of the model, it has limited representation ability. This might be one of the main reasons to explain the significant accuracy degradation for existing quantization methods when the bit-width is lower than 4.
@@
In this paper, we propose a weight quantization method, called low-rank codebook based quantization (LCQ), for LLMs. The main contributions of LCQ are outlined as follows:
@@ 2 Related Work
Existing weight quantization methods for LLMs typically employ post-training quantization (PTQ) which is relatively fast and does not require access to the full original training data (Frantar et al., 2023; Lin et al., 2024; Dettmers et al., 2023). PTQ methods such as BRECQ (Li et al., 2021), OBC (Frantar & Alistarh, 2022) and QDrop (Wei et al., 2022) are designed for quantizing traditional non-LLMs, which cannot be directly used for LLMs due to their high optimization cost. Recently, some PTQ methods have been specially designed for LLM quantization. GPTQ (Frantar et al., 2023) primarily proposes a learning method to compensate for quantization residuals and uses output reconstruction error as the objective function. However, GPTQ does not perform optimization for the quantization codebook, which results in significant accuracy loss with low bit-width. AWQ (Lin et al., 2024) and OmniQuant (Shao et al., 2024) mainly focus on learning the quantization codebook, and the quantization indices are directly computed based on the index of the quantization value that is nearest to the full-precision weight. The codebooks of AWQ and OmniQuant are rank-one, limiting the representation ability. Very recently, there appear some PTQ methods like OWQ (Lee et al., 2024), SpQR (Dettmers et al., 2024) and SquzzeLLM (Kim et al., 2024), which mainly address the issue of outliers in activations by assigning higher bit-widths to the channels/connections of activation outliers. Our LCQ is orthogonal to methods like OWQ, SpQR and SqueezeLLM, and can be further integrated with these methods.
Moreover, some LLMs compression methods (Dettmers et al., 2022; Xiao et al., 2023; Yuan et al., 2023; Wei et al., 2023) integrate weight quantization and activation quantization for further accelerating LLMs inference. However, the accuracy drops significantly when quantizing both weights and activations.
PTQ methods have also been integrated with parameter-efficient fine tuning (PEFT). Representative methods include QLoRA (Dettmers et al., 2023), PEQA (Kim et al., 2023), QA-LoRA (Xu et al., 2024), and LoftQ (Li et al., 2024). But the main goal of these methods is for PEFT rather than directly quantizing the original LLMs. Our LCQ can also be used for PEFT, but this is not the focus of this paper.
@@ 3 Methodology
As in existing weight quantization methods for LLMs (Frantar et al., 2023; Lin et al., 2024; Shao et al., 2024), we also employ post-training quantization in our LCQ. In post-training quantization, we have a calibration dataset, which is denoted as $\{\mathbf{X}_{i}\}_{i=1}^{N_{X}}$. Here, each calibration sample $\mathbf{X}_{i}\in\mathbb{R}^{L\times D}$ has a sequence length of $L$ and feature dimension of $D$, and $N_{X}$ is the total number of calibration samples. As in AWQ (Lin et al., 2024) and OmniQuant (Shao et al., 2024), we perform grouped quantization for the weights of linear layers. More specifically, for a linear layer with $N_{W}$ weights, we first divide all the weights into $N_{V}$ subsets, with each subset having $N_{W}/N_{V}$ weights. Then for each subset of weights, we group its weights into $N_{G}$ groups. We denote each subset of weights as $\mathbf{W}\in\mathbb{R}^{N_{G}\times G}$, in which $G$ denotes the weight group size and $N_{G}=\frac{N_{W}}{N_{V}\times G}$ denotes the number of weight groups per subset. A quantization codebook containing quantization values is maintained for each group of weights. For weight quantization with $b$ bits, the number of all quantization values of each group of weights is $N_{Q}=2^{b}$. We denote the quantization codebook of all groups as $\mathbf{C}\in\mathbb{R}^{N_{G}\times N_{Q}}$, the $i$th row of which is the codebook corresponding to the $i$th group of weights. The quantization function operates on the weights in an element-wise way, which quantizes each weight to the nearest quantization value in the codebook. Let $\mathbf{Z}\in[0,1,\cdots,N_{Q}]^{N_{G}\times G}$ denote the quantization index of $\mathbf{W}$, in which $Z_{i,j}$ denotes the quantization index of $W_{i,j}$. We have
@@
Then we can use the quantized weight matrix $\mathbf{W}^{Q}$ for inference. In real deployment, we only need to store the codebook matrix $\mathbf{C}$ and the quantization index matrix $\mathbf{Z}$ (Han et al., 2016). From (1), we can find that $\mathbf{Z}$ is computed based on $\mathbf{C}$. Hence, $\mathbf{C}$ is the key for quantization, which will affect accuracy, storage cost and computational cost.
@@ 3.1 Low-Rank Codebook
In AWQ (Lin et al., 2024) and GPTQ (Frantar et al., 2023), the intervals between quantization values are equal, and hence it is not necessary to explicitly store the quantization codebook $\mathbf{C}$. More specifically, $\mathbf{C}$ can be represented by two vectors, a scaling vector $\mathbf{S}_{1}\in\mathbb{R}^{1\times N_{G}}$ and a fixed quantization point set (QPS) vector $\mathbf{V}_{1}\in\mathbb{R}^{1\times N_{Q}}$. $\mathbf{V}_{1}=[-1,-1+\frac{2}{2^{b}-1},\cdots,1-\frac{2}{2^{b}-1},1]$ is a uniformly spaced vector between $[-1,1]$. $\mathbf{S}_{1}$ can be adaptively obtained based on the weight matrix $\mathbf{W}$. The codebook is then obtained by computing the outer product of $\mathbf{S}_{1}$ and $\mathbf{V}_{1}$. Furthermore, to adapt to asymmetric weight distribution, a quantization offset $\mathbf{B}\in\mathbb{R}^{N_{G}\times 1}$ is introduced, with each group of weights having its own quantization offset. The final quantization codebook can be formulated as $\mathbf{C}=\mathbf{S}^{T}_{1}\mathbf{V}_{1}-\mathbf{B}$, which is a rank-one matrix. The storage cost for $\mathbf{S}_{1}$, $\mathbf{V}_{1}$ and $\mathbf{B}$ is $O(2*N_{G}+N_{Q})$, which is much lower than $O(N_{G}*N_{Q})$ for storing $\mathbf{C}$. Although rank-one codebook can reduce the storage cost, it has limited representation ability.
@@
When $N_{D}>1$, the rank of the codebook $\mathbf{C}$ can be larger than one. When $N_{D}=1$, the codebook $\mathbf{C}$ degenerates to rank-one codebook as in existing methods (Lin et al., 2024; Frantar et al., 2023). Hence, our LCQ has a stronger representation ability than existing methods, including existing rank-one codebook based methods as special cases. We will show that LCQ with $N_{D}>1$ will achieve better accuracy than LCQ with $N_{D}=1$ in Section 4. This verifies the stronger representation ability of our low-rank codebook compared to the rank-one codebook. Figure 1 shows the low-rank codebook with $N_{D}=2$.
Figure 1: Illustration of low-rank codebook based quantization.
@@ 3.2 Objective Function
As illustrated in Figure 1, LCQ starts with a pre-trained model and tries to find the optimal $\mathbf{S}$, $\mathbf{V}$ and $\mathbf{B}$ for minimizing the loss caused by quantization. We use the output reconstruction error of a Transformer block as the objective function like the methods in (Bai et al., 2022; Li et al., 2021; Yao et al., 2022), which can make the parameters of a Transformer block be optimized simultaneously. Compared to the methods that use the output reconstruction error of a single linear layer as the objective function (Frantar et al., 2023; Lin et al., 2024), using the output reconstruction error of a Transformer block can avoid falling into local optima and prevent the accumulation of quantization error. A Transformer block in the Transformer-based models (Vaswani et al., 2017) typically contains the following linear layers: the query/key/value linear projection layers $\mathbf{W}^{qproj},\mathbf{W}^{kproj},\mathbf{W}^{vproj}$, the output projection layer $\mathbf{W}^{oproj}$ of the multi-head self-attention, and two linear layers $\mathbf{W}^{fc1}$ and $\mathbf{W}^{fc2}$ of the feed forward network. Therefore, the parameters of a Transformer block can be represented as $\mathcal{W}=\left[\mathbf{W}^{qproj},\mathbf{W}^{kproj},\mathbf{W}^{vproj},\mathbf{W}^{oproj},\mathbf{W}^{fc1},\mathbf{W}^{fc2}\right]$. Similarly, parameters $\mathcal{S},\mathcal{V},\mathcal{B}$ are used to represent all $\mathbf{S}$s, $\mathbf{V}$s, $\mathbf{B}$s within a Transformer block. We define the objective function to minimize the norm of the difference between the output with quantized weights and the output with unquantized weights in a Transformer block as follows:
$\begin{split}\min_{\mathcal{S},\mathcal{V},\mathcal{B}}&\sum_{i=1}^{N_{X}}||h(\tilde{\mathbf{X}}_{i},\mathcal{W}^{Q})-h(\mathbf{X}^{fp}_{i},\mathcal{W})||_{F}^{2}\\ &+||h(\tilde{\mathbf{X}}_{i},\mathcal{W}^{Q})-h(\tilde{\mathbf{X}}_{i},\mathcal{W})||_{F}^{2}.\end{split}$ (4)
where $h(\cdot)$ denotes the function of a Transformer block, and $N_{X}$ is the number of calibration samples. $\mathbf{X}^{fp}_{i}$ represents the feature with all previous blocks having full precision and $\tilde{\mathbf{X}}_{i}$ represents the feature with all previous blocks quantized. $\mathcal{W}^{Q}=Q(\mathcal{W},\Phi(\mathcal{S},\mathcal{V},\mathcal{B}))$ represents the quantized weights in a Transformer block. $h(\tilde{\mathbf{X}}_{i},\mathcal{W}^{Q})$ represents the output feature of the Transformer block with all blocks quantized including this block. The objective function contains two terms, one uses $h(\mathbf{X}^{fp}_{i},\mathcal{W})$ as the reconstruction target and the other uses $h(\tilde{\mathbf{X}}_{i},\mathcal{W})$. Note that the quantization algorithm does not change the model weights $\mathbf{W}$ (Shao et al., 2024), in order to avoid forgetting existing knowledge. By fixing $\mathbf{W}$, we can also reduce the optimization overhead by reducing the number of parameters to be optimized.
@@ 3.3.1 Gradient Approximation for Quantization Function
The $argmin(\cdot)$ operation in (1) will make the gradient unable to back-propagate through this function. This makes it challenging to learn the quantization codebook. We propose a gradient approximation strategy to address this issue.
As in N2UQ (Liu et al., 2022b), we rewrite the quantization function into a form that adds up multiple segments, each corresponding to the interval between two adjacent quantization values. The form is as follows:
$\begin{split}Q(W_{i,j},\mathbf{C}_{i})&=C_{i,\pi_{i}(1)}\\ &+\sum_{k=1}^{N_{Q}-1}\left(C_{i,\pi_{i}(k+1)}-C_{i,\pi_{i}(k)}\right)\\ &*\chi_{0.5}^{(k)}\left(\text{clip}\left(\frac{W_{i,j}-C_{i,\pi_{i}(k)}}{C_{i,\pi_{i}(k+1)}-C_{i,\pi_{i}(k)}},0,1\right)\right)\end{split}$ (5)